OpenStudy (anonymous):

Find the slope of the line that passes through the origin and (8, -5).

OpenStudy (mathstudent55):

Slope is difference in y-values divided by difference in x-values. The origin has coordinates (0, 0). the other point is (8, -5) The difference in y-values is: 8 - 0 = 8 The difference in x-values is: -5 - 0 = -5 The slope = 8/(-5) = -(8/5) It makes no difference whether you subtract (8 - 0)/(-5 - 0), or the other way, (0 - 8) / ( 0 - (-5)). The slope will be the same either way. What you can't do is subtract the y's in one way and the x's in the other way. That will give a wrong answer.

OpenStudy (anonymous):

alright thank you. I have another question.

OpenStudy (anonymous):

Find the value of y so that the line passing through (6, y) and (7, 5) has a slope of -3.

OpenStudy (anonymous):

how would you do that?

OpenStudy (anonymous):

@mathstudent55